1. A fountain is produced by water that flows up the tube at 0.08 m3, and then radially through two cylindrical plates before exiting to the atmosphere. Determine the velocity and pressure of the water at point A.



2. Water flows through the 30-mm-diameter ripe at 0.002 m3/s and is ejected from the 10-mm-diameter nozzle at B. Determine the velocity and pressure of the water at point A.



3. Water flows through the transition at 0.3m3/s, which causes the water to rise within the piezometer at A to a height of hA = 350 mm. Determine the height hB.


1
Expert's answer
2021-12-13T10:34:43-0500
"0.08=2\pi(0.2)(0.05)v_a\\v_a=12.7\frac{m}{s}\\
0.08=2\pi(0.4)(0.05)v_b\\v_b=6.37\frac{m}{s}\\P_a=500(6.37^2-12.7^2)=-60.8\ kPa"

The negative sign indicates that the pressure at A is a partial vacuum.